OpenStudy (anonymous):

What are the solutions to 3x^2+15=-6x? a) -1+or-i b) -1+or-2i c) -2+or-i d)-2+or-2i please answer soon

sam (.sam.):

3x^(2)+15=-6x 3x^(2)+6x+15=0 3(x^(2)+2x+5)=0 x^(2)+2x+5=0 quadratic equation x=(-b+-sqrt(b^(2)-4ac))/(2a) where ax^(2)+bx+c=0 x=(-2+-sqr((2)^(2)-4(1)(5))/2 x=(-2+-4i)/2 x=(-2+-4i)/2 ---------------------- x=-1+2i x=-1-2i ---------------------- x=-1+2i,-1-2i

OpenStudy (anonymous):

thank you so much :)